# Probing Pre-trained Language Models on Code Changes: Insights from ReDef, a High-Confidence Just-in-Time Defect Prediction Dataset

### v1

@@ Abstract.
Just-in-Time software defect prediction (JIT-SDP) plays a critical role in prioritizing risky code changes during code review and continuous integration. However, existing datasets often suffer from noisy labels and low precision in identifying bug-inducing commits. To address this, we present ReDef (Revert-based Defect dataset), a high-confidence benchmark of function-level modifications curated from 22 large-scale C/C++ projects. Defective cases are anchored by revert commits, while clean cases are validated through post-hoc history checks. Ambiguous instances are conservatively filtered out via a GPT-assisted triage process involving multiple votes and audits. This pipeline yields 3,164 defective and 10,268 clean modifications, offering substantially more reliable labels than prior existing resources. Beyond dataset construction, we provide the first systematic evaluation of how pre-trained language models (PLMs) reason about code modifications—specifically, which input encodings most effectively expose change information, and whether models genuinely capture edit semantics. We fine-tune CodeBERT, CodeT5+, and UniXcoder under five encoding strategies, and further probe their sensitivity through counterfactual perturbations that swap added/deleted blocks, invert diff polarity, or inject spurious markers. Our results show that compact diff-style encodings consistently outperform whole-function formats across all PLMs, with statistical tests confirming large, model-independent effects. However, under counterfactual tests, performance degrades little or not at all—revealing that what appears to be robustness in fact reflects reliance on superficial cues rather than true semantic understanding. These findings indicate that, unlike in snapshot-based tasks, current PLMs remain limited in their ability to genuinely comprehend code modifications.
@@ 1. Introduction
By contrast, recent code-focused pre-trained language models (PLMs) and large language models (LLMs) have demonstrated strong performance across diverse SE tasks—such as code search, summarization, repair, and vulnerability detection—precisely because they are designed to capture the meaning of code itself, not merely its statistical correlates (Guo et al., 2022; Xia et al., 2023; Ni et al., 2023). This raises a natural question for JIT-SDP: if PLMs can effectively understand static code snapshots, can they also reason about code modifications, i.e., the relationship between the before and after states? Given that PLMs have consistently shown strong effectiveness for code understanding tasks, we therefore focus our study on PLMs as the primary models for systematic evaluation (Zhang et al., 2025b; Sun et al., 2023; Zhang et al., 2025a).
To answer this question, we first construct a high-confidence, change-centric benchmark. Next, in Stage 1, we compare alternative ways of presenting modification information to PLMs, asking which input encodings yield the most effective performance. For this, we fine-tune three widely used code PLMs—CodeBERT (Feng et al., 2020), CodeT5+ (Wang et al., 2023b), and UniXcoder (Guo et al., 2022)—as backbones. However, strong results under these encodings alone do not guarantee genuine semantic reasoning, as models may still rely on superficial cues. Therefore, in Stage 2, we deliberately perturb the representation of code changes—e.g., by swapping added and deleted code blocks, inverting edit polarity, or injecting spurious markers—to test whether PLMs degrade when true semantic information is distorted. By probing how models behave under such distortions, our protocol not only measures raw accuracy but also uncovers a blind spot in current PLM research: apparent robustness may simply reflect reliance on shallow cues rather than genuine semantic reasoning about code changes.
@@
High-Confidence JIT-SDP Dataset (ReDef): We curate ReDef (Revert-based Defect dataset), a corpus of single-function changes harvested via revert anchors and refined through GPT-assisted filtering. The dataset is further validated through manual sampling to ensure benchmark reliability, and its 10K+ (3,164 defective, 10,268 clean) scale makes it well-suited for fine-tuning PLMs.
Encoding Strategies for PLMs: We conduct the first systematic comparative study of five encoding strategies for representing code modifications as PLM inputs, providing clear empirical evidence of how input representations affect change-centric defect prediction.
Counterfactual Robustness Analysis: We design counterfactual perturbation tests to probe whether PLMs truly reason about code changes beyond surface cues, revealing limitations of current models and offering guidance for future research.
Together, these contributions provide both a reliable benchmark and empirical insights into how PLMs handle code changes, informing future defect prediction research and practice. The remainder of this paper is organized as follows. Section 2 reviews background and related work. Section 3 describes our dataset construction and PLM input strategies. Section 4 outlines the experimental setup, and Section 5 reports results. Section 6 discusses threats to validity, and Section 7 concludes.
@@ 2. Background and Related Work
This section reviews prior research on software defect prediction from three perspectives. Section 2.1 outlines classical Software Defect Prediction (SDP), which predicts defect-prone modules from static snapshots using either hand-crafted metrics or learned representations. Section 2.2 covers Just-in-Time software defect prediction (JIT-SDP), which shifts the unit of analysis to commits or patches and incorporates both engineered features and recent neural approaches. It also summarizes the limitations of existing JIT-SDP datasets, such as noise introduced by SZZ-based labeling. Finally, Section 2.3 discusses advances in applying Pre-trained Language Models (PLMs) across software engineering (SE) tasks, motivating their use for change-level defect prediction.
2.1. Software Defect Prediction (SDP)
Classical SDP focuses on artifact-level prediction from static snapshots, typically at the file, class, or module granularity. Early approaches rely on hand-crafted product metrics (e.g., size, complexity, coupling) and process/organization metrics (e.g., churn, revisions, developer ownership, past defects) (Malhotra, 2015; Bell et al., 2013; Huda et al., 2017). These features are commonly fed into traditional learners such as logistic regression, random forests, and Support Vector Machines (SVMs). For example, Elish et al. showed that SVMs perform competitively on NASA datasets compared to logistic regression (Elish and Elish, 2008), while ensemble methods based on random forests also address class imbalance (Balaram and Vasundra, 2022). Decision tree–based models have similarly achieved strong performance (Wang et al., 2012; Khoshgoftaar and Seliya, 2002). More recent work replaces hand-crafted metrics with learned representations, leveraging topic models (Nguyen et al., 2011), deep generative models (Wang et al., 2018), or transformer-based encoders (Omri and Sinz, 2020). Evaluation often considers effort-aware measures (e.g., recall at top-x% LOC) and cross-project transferability (Bell et al., 2013; Ostrand and Weyuker, 2007; Tosun et al., 2009; Turhan et al., 2009). Despite this progress, snapshot-based prediction cannot capture the semantics of specific changes, and file- or module-level labels blur defective regions—complicating causal interpretation and limiting actionable insights for change-level prediction.
2.2. Just-in-Time Software Defect Prediction (JIT-SDP) and Datasets
@@
More recently, researchers have applied PLMs to JIT-SDP. Guo et al. systematically compared six transformer-based backbones (RoBERTa, CodeBERT, BART, PLBART, GPT2, and CodeGPT) and showed that PLM-based variants consistently outperform traditional baselines such as DeepJIT and CC2Vec (Guo et al., 2023). Building on this, Abu Talib et al. demonstrated that parameter-efficient fine-tuning (PEFT) techniques such as LoRA can further boost JIT-SDP performance on ApacheJIT (Abu Talib et al., 2024), while Lin et al. introduced CCT5—a CodeT5-initialized model pre-trained on CodeChangeNet—that achieves state-of-the-art results on code-change tasks including JIT-SDP (Lin et al., 2023). Jiang et al. extended CodeBERT with a Replaced Message Identification objective, resulting in BiCC-BERT and JIT-BiCC, which improved F1-scores by over 10% (Jiang et al., 2025). Together, these studies highlight the promise of PLMs for JIT-SDP, yet it remains unclear whether current gains reflect genuine modeling of code-change semantics or incidental signals from existing datasets—underscoring the need for both more reliable benchmarks and systematic evaluation of model behavior.
Despite these advances, a critical limitation of existing JIT-SDP datasets lies in label reliability. Afric et al. reported that even with RoBERTa-based issue classification, file-level datasets contained on average 14.3% mislabeled instances (Afric et al., 2023). Herbold et al. showed that SZZ-based labeling is particularly error-prone, with only about half of the commits identified as bug-inducing being correct (Herbold et al., 2022). Several variants of SZZ attempt to mitigate this noise—for example, by excluding refactoring edits that standard SZZ mistakenly flags as defect-inducing or by tracing commit chains more accurately to recover missed bug-introducing links (Niu et al., 2025; Lyu et al., 2024). However, even with such refinements, existing datasets still fall short in reliability, underscoring the need for explicit validation of defect labels, which we address in RQ1.
2.3. Pre-trained Language Models (PLMs) in Other Software Engineering Tasks
Pre-trained language models (PLMs) trained on large code corpora—such as CodeBERT, GraphCodeBERT, UniXCoder, and CodeT5—have been increasingly applied to diverse SE tasks, often achieving state-of-the-art results.
In code review, Li et al. proposed CodeReviewer, a Transformer-based encoder–decoder PLM tailored for diff review, improving both comment generation and code refinement quality (Li et al., 2022). Similarly, Kartal et al. demonstrated that transformer-based similarity measures outperform prior baselines in automated comment recommendation (Kartal et al., 2024).
In program repair, PLMs have achieved significant progress. For instance, Mashhadi and Hemmati applied CodeBERT to repair Java bugs, achieving 19–72% fix accuracy on ManySStuBs4J datasets (Mashhadi and Hemmati, 2021). Retrieval-augmented methods like RAP-Gen combine CodeT5 with external bug-fix retrieval, showing marked improvements over previous approaches—repairing 15 more bugs on the Defects4J benchmark (Wang et al., 2023a). Additionally, AlphaRepair uses CodeBERT in a zero-shot, infilling-style repair setup and achieves state-of-the-art results without task-specific retraining (Xia et al., 2023).
In vulnerability detection, PLMs have shown strong effectiveness across diverse settings. For example, one study fused multi-modal knowledge (structural, semantic, and syntactic features) with PLM encoders to achieve state-of-the-art function-level vulnerability detection (Ni et al., 2023). Atiiq et al. fine-tuned PLMs on the multi-language CVEFixes dataset (JavaScript, Java, Python, PHP, Go, C/C++), reporting substantial variation across languages and finding that UniXcoder outperformed other models (Atiiq et al., 2024). More recently, a comprehensive benchmark demonstrated that PLMs can surpass much larger code LLMs in accuracy and F1-scores, underscoring their superior inductive bias and efficiency for vulnerability detection tasks (Zhang et al., 2025b).
Taken together, PLMs have shown impressive capabilities across diverse SE tasks. Importantly, many of these tasks—such as program repair and vulnerability detection—require reasoning about the semantics of code edits or risky patterns, which parallels the demands of defect prediction. Their strong performance in such semantics-sensitive settings suggests that PLMs are particularly promising for change-aware defect prediction, motivating our systematic investigation in this paper.
@@ 3. Methodology
Our methodology comprises two main pillars. (A) ReDef Construction: We curate defective and clean code modifications from large-scale C/C++ projects, leveraging revert signals, post-hoc history screening, and expert audits to ensure high label reliability. (B) PLM Evaluation Protocol: We design a two-stage framework to assess how models interpret code modifications. Figure 1 summarizes the overall approach.
@@ 3.1. ReDef Construction
Defective Modification. We extracted commits that touched exactly one file and whose modified lines were confined to a single function. This design ensures reliable causal attribution to a specific function—a choice also adopted in recent datasets such as PrimeVul (Ding et al., 2024). From this single-function pool, we identified defective modifications by collecting revert commits whose messages (i) contained the keyword “revert” and (ii) explicitly referenced the hash of an earlier commit. Since such reverts typically indicate that the referenced commit introduced a problem, the targeted commits are highly likely to represent defective modification. For each referenced hash, we retrieved its associated commit (the target of the revert) and collected four artifacts for analysis: the reverted commit message, the function_before and function_after from that commit, and the revert commit message itself. Together these formed an evidence bundle and were submitted to GPT-4o three independent times under a fixed rubric to remove ambiguous or non-defect reverts (e.g., formatting changes, policy reversions, or hash-only messages without rationale). The rubric distinguished each revert as: (A) Failed/incomplete fix (DefectiveDefective), (B) Bug introduction (CleanDefective), or (C) Other (non-defect). Categories (A) and (B) were merged into the defective class, while (C) served as a filter to discard non-defect cases. To ensure high precision, a candidate was retained only if all three GPT-4o votes unanimously fell in {A,B} (e.g., A/A/A or A/B/B). Any sample with one or more (C) votes (e.g., A/A/C) was conservatively discarded.
@@
After assembling the labels, each repository was split time-respectively (80%/10%/10% into train/validation/test by commit timestamp) to avoid temporal leakage—training on commits occurring after the test set would otherwise expose future information. Class balance (defective/clean) was preserved within each split, and splitting was performed per repository to ensure that every project was represented across partitions, preventing cases where a project might be entirely absent from validation or test.The per-project partitions were then concatenated to form the final splits.
To characterize input length, we tokenized function_after using each of the PLMs employed in our experiments (CodeBERT, UniXcoder, CodeT5+). As shown in Figure 3, the cumulative distribution indicates that although the exact counts vary slightly by model, around 55–65% of functions exceed the 512-token limit, meaning the entire function cannot be fed to the model without truncation. With UniXcoder’s 1,024-token limit, roughly 70% of functions fit in full. This distribution motivates a comparison between encoding strategies that provide only the edited lines versus those that supplying the full function at the risk of truncation. Although the dataset is derived from commit history, the intended inference unit is a function-level modification, which can be evaluated pre-commit (e.g., in an Integrated Development Environment (IDE) or local CI hooks) as well as at review/merge time. To assess label reliability, two authors independently audited a stratified sample of 100 defective commits under a shared rubric and adjudicated disagreements.
3.2. PLM Evaluation Protocol
Our goal is to assess whether pre-trained language models (PLMs) can effectively leverage code-change information for defect prediction. To this end, we designed a two-stage evaluation. Stage 1 compares alternative input encodings that vary in how much context and change signal are provided. Stage 2 introduces counterfactual perturbations that deliberately distort change signals, checking whether models degrade as expected when edit semantics are disrupted. Together, these stages provide a structured protocol for evaluating whether PLMs rely on genuine semantic cues rather than superficial patterns.
In Stage 1, we compared five alternative input encodings, each tokenized using the model’s native tokenizer before being fed into the PLM encoder.

### v2

# ReDef: Do Code Language Models Truly Understand Code Changes for Just-in-Time Software Defect Prediction?

@@ Abstract.
Just-in-Time software defect prediction (JIT-SDP) plays a critical role in prioritizing risky code changes during code review and continuous integration. However, existing datasets often suffer from noisy labels and low precision in identifying bug-inducing commits. To address this, we present ReDef (Revert-based Defect dataset), a high-confidence benchmark of function-level modifications curated from 22 large-scale C/C++ projects. Defective cases are anchored by revert commits, while clean cases are validated through post-hoc history checks. Ambiguous instances are conservatively filtered out via a GPT-assisted triage process involving multiple votes and audits. This pipeline yields 3,164 defective and 10,268 clean modifications, offering substantially more reliable labels than prior resources. Beyond dataset construction, we provide a systematic evaluation of how Code Language Models (CLMs)—specifically CodeBERT, CodeT5+, UniXcoder, and Qwen2.5—reason about code modifications. We first investigate which input encodings most effectively expose change information under five different strategies. We then design four counterfactual perturbation strategies (e.g., swapping added/deleted blocks, inverting diff polarity) to serve as diagnostic probes. We posit that if models genuinely capture change semantics, such distortions should lead to a clear decline in predictive performance. Our results show that compact diff-style encodings consistently outperform whole-function formats across all CLMs, supported by rigorous statistical confirmation. However, under counterfactual tests, performance remains effectively stable, revealing that what appears to be robustness in fact reflects a reliance on superficial cues rather than true semantic understanding. These findings indicate that, at least in code-change understanding tasks, current CLMs remain limited in their ability to genuinely comprehend the relational dynamics of code modifications.
@@ 1. Introduction
By contrast, recent Code Language Models (CLMs)—ranging from specialized encoder-based architectures to large-scale generative decoders (LLMs)—have demonstrated an unprecedented ability to capture the deep semantic meaning of code, moving beyond mere statistical correlates. This advancement has led to state-of-the-art performance across diverse software engineering tasks, including code search, summarization, repair, and vulnerability detection (Guo et al., 2022; Xia et al., 2023; Ni et al., 2023; Hui et al., 2024). This success raises a fundamental question for JIT-SDP: if CLMs can effectively understand static code snapshots, can they also reason about code modifications (i.e., the causal relationship between the before and after states)? Given that CLMs, regardless of their specific architecture, have consistently shown superior effectiveness in code understanding, we focus our study on evaluating these models as the primary candidates for a systematic evaluation of change-level defect prediction (Zhang et al., 2025b; Sun et al., 2023; Zhang et al., 2025a).
To answer this question, we first construct a high-confidence, change-centric benchmark. This high label precision is essential for a rigorous analysis, as it ensures that any observed model limitations stem from a lack of semantic understanding rather than noise in the ground truth. Next, in Stage 1, we compare alternative ways of presenting modification information to CLMs, asking which input encodings yield the most effective performance. For this, we fine-tune four widely used CLMs—CodeBERT (Feng et al., 2020), CodeT5+ (Wang et al., 2023), UniXcoder (Guo et al., 2022), and Qwen2.5 (Hui et al., 2024)—as backbones. However, strong results under these encodings alone do not guarantee genuine semantic reasoning, as models may still rely on superficial cues. Therefore, in Stage 2, we deliberately perturb the representation of code changes—e.g., by swapping added and deleted code blocks, inverting edit polarity, or injecting spurious markers—to test whether CLMs degrade when true semantic information is distorted. By probing how models behave under such distortions, our protocol not only measures raw accuracy but also uncovers a blind spot in current CLM research: apparent robustness may simply reflect reliance on shallow cues rather than genuine semantic reasoning about code changes.
@@
High-Confidence JIT-SDP Dataset (ReDef): We curate ReDef (Revert-based Defect dataset), a corpus of single-function changes harvested via revert anchors and refined through GPT-assisted filtering. The dataset is further validated through manual sampling to ensure benchmark reliability, and its 10K+ (3,164 defective, 10,268 clean) scale makes it well-suited for fine-tuning CLMs.
Encoding Strategies for CLMs: We conduct the first systematic comparative study of five encoding strategies for representing code modifications as CLM inputs, providing clear empirical evidence of how input representations affect change-centric defect prediction.
Counterfactual Robustness Analysis: We design counterfactual perturbation tests to probe whether CLMs truly reason about code changes beyond surface cues, revealing limitations of current models and offering guidance for future research.
Together, these contributions provide both a reliable benchmark and empirical insights into how CLMs handle code changes, informing future defect prediction research and practice. The remainder of this paper is organized as follows. Section 2 reviews background and related work. Section 3 describes our dataset construction and two-stage evaluation protocol, encompassing both input encoding strategies and counterfactual perturbation methods. Section 4 outlines the experimental setup, and Section 5 reports results. Section 6 discusses threats to validity, and Section 7 concludes.
@@ 2. Background and Related Work
This section reviews prior research on software defect prediction and the challenges of dataset reliability. Section 2.1 briefly outlines classical Software Defect Prediction (SDP), which focuses on identifying defect-prone modules from static snapshots. Section 2.2 covers the evolution of Just-in-Time software defect prediction (JIT-SDP), focusing on the transition from hand-crafted metrics to deep learning and recent Code Language Models (CLMs). Finally, Section 2.3 discusses the critical limitations of existing JIT-SDP datasets, specifically the noise and reliability issues inherent in SZZ-based labeling, which motivates our construction of the ReDef benchmark.
2.1. Artifact-level Software Defect Prediction (SDP)
Classical SDP focuses on predicting defect-prone artifacts (e.g., files or modules) from static snapshots. Early approaches relied on hand-crafted product metrics (e.g., size, complexity, coupling) and process metrics (e.g., churn, developer ownership, past defects) (Malhotra, 2015; Bell et al., 2013; Huda et al., 2017) fed into traditional machine learning models like SVMs and random forests (Elish and Elish, 2008; Balaram and Vasundra, 2022; Wang et al., 2012; Khoshgoftaar and Seliya, 2002). More recent work has shifted toward learned representations using deep generative models or transformer-based encoders to capture richer features (Wang et al., 2018; Omri and Sinz, 2020). While these methods have evolved to consider various practical evaluation settings (Ostrand and Weyuker, 2007; Tosun et al., 2009), snapshot-based prediction remains limited as it cannot capture the semantics of specific code changes and often blurs defective regions at a coarse granularity. This lack of change-level precision necessitates Just-in-Time (JIT) approaches that provide actionable insights at the time of commit.
2.2. Just-in-Time Software Defect Prediction (JIT-SDP)
@@
More recently, researchers have applied CLMs to JIT-SDP. Guo et al. systematically compared six transformer-based backbones (RoBERTa, CodeBERT, BART, PLBART, GPT2, and CodeGPT) and showed that CLM-based variants consistently outperform traditional baselines such as DeepJIT and CC2Vec (Guo et al., 2023). Building on this, Abu Talib et al. demonstrated that parameter-efficient fine-tuning (PEFT) techniques such as LoRA can further boost JIT-SDP performance on ApacheJIT (Abu Talib et al., 2024), while Lin et al. introduced CCT5—a CodeT5-initialized model pre-trained on CodeChangeNet—that achieves state-of-the-art results on code-change tasks including JIT-SDP (Lin et al., 2023). Jiang et al. extended CodeBERT with a Replaced Message Identification objective, resulting in BiCC-BERT and JIT-BiCC, which improved F1-scores by over 10% (Jiang et al., 2025). Together, these studies highlight the promise of CLMs for JIT-SDP, appearing to possess an advanced capability to comprehend code modifications. However, most existing research has focused on compact encoder-based or encoder-decoder models, leaving the potential of contemporary large-scale decoder-based models—such as the Qwen (Yang et al., 2025) or Llama families (Grattafiori et al., 2024)—largely underexplored in this domain. Furthermore, it remains unclear whether these performance gains reflect genuine semantic understanding or a reliance on superficial textual patterns, underscoring the need for a systematic evaluation of model behavior.
2.3. Labeling Reliability and SZZ Limitations
A precise and high-confidence dataset is a fundamental prerequisite for evaluating whether Code Language Models truly understand the semantics of code changes. Despite the technical progress in neural modeling described in Section 2.2, a critical bottleneck in JIT-SDP remains the reliability of dataset labeling. Most existing JIT-SDP datasets rely on SZZ-based labeling, which identifies bug-introducing commits by tracing the origins of the code lines modified in bug-fixing commits (Śliwerski et al., 2005). However, this approach introduces substantial noise that propagates labeling errors and obscures evaluation quality. Herbold et al. showed that SZZ-based labeling is highly error-prone, with only about half of the commits identified as bug-inducing being correct (Herbold et al., 2022). Even after additional refinement using a RoBERTa-based issue classifier, Afric et al. reported that datasets still contained, on average, 14.3% mislabeled instances (Afric et al., 2023). Lyu et al. conducted a large-scale evaluation on over 76K developer-labeled Linux kernel commits and showed that SZZ variants suffer from a recall degradation of 13.8%, with 17.47% of bug-fixing commits being ghost commits (Lyu et al., 2024).
To mitigate these errors, several SZZ variants have been proposed (Fan et al., 2019). Early improvements like Annotation Graph SZZ (AG-SZZ) attempted to filter non-semantic changes (e.g., whitespace and comments) (Kim et al., 2006). Subsequent variants addressed meta-changes (MA-SZZ) (Da Costa et al., 2016) and automated refactoring detection (RA-SZZ) (Neto et al., 2018) to prevent incorrect blame assignment. More recent efforts include TC-SZZ, which iteratively traces deeper histories to recover missed commits (Lyu et al., 2024), and Code Change Tactics (CAT), which disentangles refactoring-propagated changes from actual defect-inducing edits (Niu et al., 2025). Yet, even with these sophisticated refinements, automated labeling still struggles with the inherent ambiguity of commit histories.
Fundamentally, pinpointing the root cause of a defect requires deep domain expertise; without explicit documentation in commit messages, it is notoriously difficult for external researchers—or automated tools—to accurately assign blame to specific code modifications. Furthermore, even manual inspection by non-developers remains subjective and error-prone, as identifying a bug’s origin from raw diffs alone lacks necessary semantic grounding. These persistent reliability gaps underscore the need for more conservative and evidence-based labeling strategies that rely on explicit developer feedback rather than speculative history tracking. In this paper, we address this need by introducing the ReDef benchmark, which prioritizes high-confidence signals to provide a more robust environment for assessing the true reasoning capabilities of various Code Language Models, spanning encoder-based models (CodeBERT, UniXcoder, and the encoder of CodeT5+) and a large-scale decoder-based model (Qwen2.5).
@@ 3. Methodology
Our methodology comprises two main pillars. (A) ReDef Construction: We curate defective and clean code modifications from large-scale C/C++ projects, leveraging revert signals, post-hoc history screening, and expert audits to ensure high label reliability. (B) Code Language Models Evaluation Protocol: We design a two-stage framework to assess how various Code Language Models (CLM) interpret code modifications. Figure 1 summarizes the overall approach.
@@ 3.1. ReDef Construction
Defective Modification. We extracted commits that touched exactly one file and whose modified lines were confined to a single function. This design ensures reliable causal attribution to a specific function—a choice also adopted in recent datasets such as PrimeVul (Ding et al., 2024). From this single-function pool, we identified defective modifications by collecting revert commits whose messages (i) contained the keyword “revert” and (ii) explicitly referenced the hash of an earlier commit. Since such reverts typically indicate that the referenced commit introduced a problem, the targeted commits are highly likely to represent defective modifications.
For each referenced hash, we retrieved its associated commit (the target of the revert) and collected four artifacts for analysis: the reverted commit message, the function_before and function_after from that commit, and the revert commit message itself. Together these formed an evidence bundle and were submitted to GPT-4o three independent times under a fixed rubric to remove ambiguous or non-defect reverts (e.g., formatting changes, policy reversions, or hash-only messages without rationale). The rubric distinguished each revert as: (A) Failed/incomplete fix (DefectiveDefective), (B) Bug introduction (CleanDefective), or (C) Other (non-defect). Categories (A) and (B) were merged into the defective class, while (C) served as a filter to discard non-defect cases. To ensure high precision, a candidate was retained only if all three GPT-4o votes unanimously fell in {A,B} (e.g., A/A/A or A/B/B). Any sample with one or more (C) votes (e.g., A/A/C) was conservatively discarded.
Across the 22 subject projects, we identified approximately 0.43M reverts out of 3.37M fix-related commits, indicating that reverts account for roughly 13% of the total defect pool. This proportion further decreases after applying the single-function constraint and GPT-4o triage. While this design inevitably limits coverage, it reflects a deliberate trade-off to prioritize label precision and mitigate the “falsely blaming” issue inherent in traditional SZZ-based heuristics. In contrast to heuristic lineage tracking, revert commits provide explicit causal links between bug fixes and their corresponding bug-inducing changes, and our filtering pipeline removes tangled or ambiguous cases. As a result, ReDef complements existing larger but noisier datasets by offering a controlled, high-confidence benchmark for evaluating the semantic reasoning capabilities of code models.
@@
Data Partitioning and Statistics. After assembling the labels, each repository was split time-respectively (80%/10%/10% into train/validation/test by commit timestamp) to avoid temporal leakage—training on commits occurring after the test set would otherwise expose future information. Class balance (defective/clean) was preserved within each split, and splitting was performed per repository to ensure that every project was represented across partitions, preventing cases where a project might be entirely absent from validation or test. The per-project partitions were then concatenated to form the final splits. Table 2 summarizes the scale of the curated modifications. On average, the studied functions consist of 94.52 Lines of Code (LOC), with modifications typically affecting 22.82% of the code. The gradual increase in average LOC from training to test sets reflects the natural growth of project complexity over time, as our chronological split captures the inherent evolution of software projects. Conversely, the slight decrease in the percentage of changed code suggests that modifications become more localized as projects mature and stabilize.
| Split | Count | Avg. LOC | % Changed |
| Train | 10,744 | 92.12 | 23.30% |
| Valid | 1,344 | 102.29 | 22.01% |
| Test | 1,344 | 105.98 | 19.81% |
| Overall | 13,432 | 94.52 | 22.82% |
To characterize input length, we tokenized function_after using each of the CLMs employed in our experiments (CodeBERT, UniXcoder, CodeT5+, Qwen2.5). As shown in Figure 3, the cumulative distribution indicates that although the exact counts vary slightly by model, around 48–65% of functions exceed the 512-token limit, meaning the entire function cannot be fed to the model without truncation. With Qwen2.5’s 1,024-token limit, roughly 77% of functions fit in full. This distribution motivates a comparison between encoding strategies that provide only the edited lines versus those that supplying the full function at the risk of truncation. Although the dataset is derived from commit history, the intended inference unit is a function-level modification, which can be evaluated pre-commit (e.g., in an Integrated Development Environment (IDE) or local CI hooks) as well as at review/merge time. To assess label reliability, two authors independently audited a stratified sample of 100 defective commits under a shared rubric and adjudicated disagreements.
3.2. Code Language Models Evaluation Protocol
Our goal is to assess whether Code Language Models (CLMs) can effectively leverage code-change information for defect prediction. To this end, we designed a two-stage evaluation. Stage 1 compares alternative input encodings that vary in how much context and change signal are provided. Stage 2 introduces counterfactual perturbations that deliberately distort change signals, checking whether models degrade as expected when edit semantics are disrupted. Together, these stages provide a structured protocol for evaluating whether CLMs rely on genuine semantic cues rather than superficial patterns.
In Stage 1, we compared five alternative input encodings, each tokenized using the model’s native tokenizer before being fed into the CLM encoder.